dx/(x^2-y z) =dy/(y^2-zx)=dz/(z^2-xy)
1
Expert's answer
2021-10-01T13:18:39-0400
"\dfrac{dx }{x^2-yz}=\dfrac{dy}{y^2-zx}=\dfrac{dz}{z^2-xy}"

Taking the multipliers "y, z, x," then the ratio,

"\dfrac{ydx+zdy+xdz }{x^2y-y^2z+y^2z-z^2x+xz^2-x^2y}"

"=\dfrac{ydx+zdy+xdz }{0}"

Hence


"ydx+zdy+xdz=0"

Integrate


"xy+yz+zx=c_1"

Taking the multipliers "xy, xz, x^2," then the ratio,

"\dfrac{xydx+xzdy+x^2dz }{x^3y-xy^2z+xy^2z-z^2x^2+x^2z^2-x^3y}"

"=\dfrac{xydx+xzdy+x^2dz }{0}"

Hence


"xydx+xzdy+x^2dz=0"

Integrate


"\dfrac{x^2y}{2}+xyz+x^2z=\dfrac{c_2}{2}"

"x^2y+2xyz+2x^2z=c_2"

Therefore the solutin is


"\Phi(xy+yz+zx,x^2y+2xyz+2x^2z )=0"

where "\Phi" is an arbitrary function.